A metal strip 2.50 cm wide and 1.80 cm thick carries a current of 25.9 A in a region containing a perpendicular uniform magnetic field of strength 1.65 T. The Hall voltage across the 2.50 cm width of the strip is measured to be 4.36 μV. Calculate the drift speed of the electrons in the strip.

Find the number density of the charge carriers in the strip (in m-3).
Answer #1

A)

Drift velocity, Vd = (V/w)/B

Vd = (4.36 x 10^-6)/(1.65 x 0.025) = 1.06 x 10^-4 m/s

B)

Number density, n = (I/wt) /e Vd

n = (25.9)(0.025 x 0.018 x 1.6 x 10^-19 x 1.06 x 10^-4) = 3.40 x 10^27 /m^3

Comment in case any doubt please rate my answer.....